Algebra

derivation of volume ofcone

anandkumar
12 Years agoGrade

1 Answer

Sher Mohammad
12 Years ago
it can be drived using the fact volume of cone is 1/3rd of cylinder,
base area of cylinder= pi*r^2
cylinder volume=area* h=pi*r^2*h
cone volume =1/3cylinder volume=pi*r^2*h/3

sher mohammad
b.tech, iit delhi